OpenStudy (anonymous):

Use differentials to approximate the error and relative error in using the given measurement to calculate the value of the function, if the measurement may be off by the specified amount. (a) Using radius 10 miles for the area of a circular island if the radius may vary by 0.02 miles -approximate error: -approximate relative error: (b) Using volume V = 4 m^3 for the pressure P = 220/V of a gas in a balloon if the volume may be off by 0.2 m^3 -approximate error: -approximate relative error: (c) Using an intensity of I = 64 lumens for the distance -approximate error: -approximate

OpenStudy (phi):

how about area A= \( \pi r^2 \) d A= \( 2\pi r \ dr \) now replace dr with 0.02 and r with 10 to find d A

OpenStudy (anonymous):

ok sec

OpenStudy (anonymous):

2pi(10)(0.02) = 1.25664

OpenStudy (anonymous):

@phi what do i do now?

OpenStudy (phi):

I think the relative error is error divided by area

OpenStudy (anonymous):

oh ok let me try that!

OpenStudy (phi):

You may want to write the relative error as a percent

OpenStudy (anonymous):

yay that worked!

OpenStudy (anonymous):

@phi thanks soo much... Would u help me with part b too?

OpenStudy (phi):

can you take the derivative of P = 220/V ?

OpenStudy (anonymous):

-220/v^2?

OpenStudy (phi):

you should also show the differentials. in other words derivative of P is dP and the derivative of v^-1 is -V^-2 dV so you get \[ dP = - \frac{220}{V^2} dV \] now find dP using the data they gave you.

OpenStudy (anonymous):

i get -2.75 but it says is wrong

OpenStudy (anonymous):

(-220/4^2)*0.2=-2.75 but it says the answer is wrong for the approximate error

OpenStudy (phi):

they may not want the minus sign. they don't say the error is plus or minus

OpenStudy (anonymous):

yup that worked!

OpenStudy (anonymous):

so now i do the voluma divided by dP?

OpenStudy (phi):

relative error: error/"truth"

OpenStudy (phi):

so the other way round dP/P

OpenStudy (anonymous):

sweet thank you soo much!!